# Low-Altitude ISAC with Rotatable Active and Passive Arrays

### v1

@@ II-B Direction and Steering Vectors
For a far-field direction parameterized by an elevation-azimuth pair $(\vartheta,\varphi)$ with elevation from the $xOy$ plane and azimuth from the $x$-axis, define the direction vector
@@
Accordingly, we use the following indexed notational shorthands for distinct BS-user (BU), RIS-BS (RB), BS-RIS (BR), and RIS-user (RU) links, respectively
@@
$\!\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{BR}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{BR}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}\!,\!\!$ (7c)
$\!\boldsymbol{t}_{k,q}^{\text{RU}}(\boldsymbol{r}^{\text{B}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,q}^{\text{RU}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,q}^{\text{RU}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}\!.\!\!$ (7d)
@@
$\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{B}})=\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{RT}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{RT}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}.$ (9b)
@@ II-C Rotation-Aware Multipath Channels
$\boldsymbol{n}(\boldsymbol{r}^{\mathrm{A}})=\boldsymbol{R}(\boldsymbol{r}^{\mathrm{A}})\bar{\boldsymbol{n}}$ (10)
@@
where $G_{0}^{\text{B}}$ and $G_{0}^{\text{R}}$ denotes the maximum gain achieved when the signal aligns with the boresight, and $p^{\text{B}}$ and $p^{\text{R}}$ are the directivity factors controlling the main-lobe width, following from directivity normalization over the full sphere. Then, with given numbers of resolvable paths $L_{k}$, $P$, and $Q_{k}$, and small-scale complex gains $a_{k,l}$, $b_{p}$, and $c_{k,q}$ that absorb path loss and fading, the BS-to-user channel $\boldsymbol{h}_{k}\left(\boldsymbol{r}^{\text{B}}\right)\in\mathbb{C}^{M}$, BS-to-RIS channel $\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\in\mathbb{C}^{M\times N}$, and RIS-to-user channel $\boldsymbol{g}_{k}\left(\boldsymbol{r}^{\text{R}}\right)\in\mathbb{C}^{N}$ are respectively expressed as
$\boldsymbol{h}_{k}\left(\boldsymbol{r}^{\text{B}}\right)=\sum_{l=1}^{L_{k}}{a_{k,l}\sqrt{G_{k,l}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}}\boldsymbol{t}_{k,l}^{\text{BU}}\left(\boldsymbol{r}^{\text{B}}\right)$ (12a)
$\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)=\sum_{p=1}^{P}{b_{p}\sqrt{G_{k,p}^{\text{B}}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)G_{k,p}^{\text{R}}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}}$
$\qquad\qquad\qquad\quad\,\,\times\boldsymbol{t}_{p}^{\text{BR}}\left(\boldsymbol{r}^{\text{B}}\right)\left(\boldsymbol{t}_{p}^{\text{RB}}\left(\boldsymbol{r}^{\text{R}}\right)\right)^{H}$ (12b)
$\boldsymbol{g}_{k}\left(\boldsymbol{r}^{\text{R}}\right)=\sum_{q=1}^{Q_{k}}{c_{k,q}\sqrt{G_{k,q}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}}\boldsymbol{t}_{k,l}^{\text{RU}}\left(\boldsymbol{r}^{\text{R}}\right).$ (12c)
@@
$\boldsymbol{f}_{k}(\boldsymbol{r}^{\text{B}}\!,\boldsymbol{r}^{\text{R}}\!,\boldsymbol{\theta})\!=\!\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})\!+\!\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\text{diag}(\boldsymbol{\theta})\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}}),\!\!$ (13)
@@
$\boldsymbol{f}_{\text{S},a}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})=$ $\sqrt{G_{a}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}a_{a}\boldsymbol{t}_{a}^{\text{BT}}\left(\boldsymbol{r}^{\text{B}}\right)+\!\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\text{diag}(\boldsymbol{\theta})$
$\times\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{B}}\right)}c_{a}\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{B}})$ (14)
II-D Transmit Signal, Communication SINR, and Sensing MSE
@@
where $\boldsymbol{W}_{\text{C}}=[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{K}\boldsymbol{]}\in\mathbb{C}^{M\times K}$ collects $K$ communication beams, $\boldsymbol{W}_{\text{S}}\in\mathbb{C}^{M\times M}$ contains dedicated sensing beams, $\boldsymbol{s}_{\text{C}}=[s_{1},\dots,s_{K}]^{T}$ denotes the communication symbols, and $\boldsymbol{s}_{\text{S}}$ denotes the sensing signals, with $\mathbb{E}\left[\boldsymbol{s}_{\text{S}}\boldsymbol{s}_{\text{S}}^{H}\right]=\boldsymbol{I}_{M}$, $\boldsymbol{s}_{\text{C}}\sim\mathcal{C}\mathcal{N}\left(\boldsymbol{0},\boldsymbol{I}_{K}\right)$, and $\mathbb{E}\left[\boldsymbol{s}_{\text{S}}\boldsymbol{s}_{\text{C}}^{H}\right]=0$. Let $\boldsymbol{W}\triangleq\left[\boldsymbol{W}_{\text{C}},\boldsymbol{W}_{\text{S}}\right]$ denotes the BS ISAC transmit beamforming matrix, $\boldsymbol{s}\triangleq\left[\boldsymbol{s}_{\text{C}}^{T},\boldsymbol{s}_{\text{S}}^{T}\right]^{T}$, the received signal at the communication user $k$ is written as
$y_{k}$ $=\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{Ws}$ (16a)
$=\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{w}_{k}s_{k}+\sum_{i=1,i\neq k}^{K}{\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{w}_{i}}x_{i}$
@@
where $n_{k}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{k}^{2}\right)$ is the complex additive white Gaussian noise. Taking further $\boldsymbol{W}\triangleq\left[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{K+N}\right]$, where $\boldsymbol{w}_{i}$ is the $i$-th column of $\boldsymbol{W}$, the SINR of the $k$-th user is
@@
Given a discretized grid $\{(\vartheta_{a},\varphi_{a})\}_{a=1}^{A},a\in\{1,\dots,A\}$ with $A$ sampled points in total, which samples the azimuth and elevation angular domain over the target sensing area $\mathcal{A}$, along with the desired beam pattern $\mathcal{P}_{\text{d}}(\vartheta_{a},\varphi_{a})$, the mean squared error (MSE) between $\mathcal{P}$ and $\mathcal{P}_{\text{d}}$ to evaluate the similarity of the sensing beam pattern is defined as []
@@
A typical ideal desired beam pattern is given by
@@
where $\mathcal{T}$ is the set of focused grid points in the region $\mathcal{A}$.
@@ II-E Problem Formulation
Accordingly, we jointly optimize the BS and RIS rotations, the RIS phase shifts, and the BS transmit beamforming to maximize the sum-rate performance for communication with the guaranteed sensing requirement on the beam pattern MSE
$\text{(P1)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\right)}$ (22a)
s.t. $\text{MSE}\leq\eta_{\text{th}}$ (22b)
$\|\boldsymbol{W}\|_{F}^{2}=\text{Tr}(\boldsymbol{WW}^{H})\leq P_{\text{B}},$ (22c)
$|\theta_{n}|=1,\forall n\in\mathcal{N},$ (22d)
$\zeta_{i}^{\text{B}}\!\in\left[\zeta_{i,\min}^{\text{B}},\zeta_{i,\max}^{\text{B}}\right],i\in\{x,y,z\},$ (22e)
$\zeta_{i}^{\text{R}}\!\in\left[\zeta_{i,\min}^{\text{R}},\zeta_{i,\max}^{\text{R}}\right],i\in\{x,y,z\},$ (22f)
where $\eta_{\text{th}}$ in (22b) is the maximum allowed sensing MSE, $P_{\mathrm{B}}$ in (22c) is the BS power budget, the unit-modulus constraint (23d) enforces a passive RIS, and the box constraints (22e) and (22f) capture mechanical limits of the array rotations at BS and RIS, respectively.
From the formulated problem, we observe that the rotation variables $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ provide a geometric control that affects the steering vectors and directive gain, thereby complementing the electromagnetic control offered by active and passive beamforming at the RIS and BS arrays; consequently, jointly optimizing these blocks yields a significant DoF for low-altitude ISAC. However, (P1) is highly non-convex due to unit-modulus constraints, rotation-induced nonlinearities in the steering vector and directive gain, and the coupling among $\boldsymbol{r}^{\text{B}}$, $\boldsymbol{r}^{\text{R}}$, $\boldsymbol{\theta}$, and $\boldsymbol{W}$ in both SINR and MSE expressions.

### v2

@@ II-B Direction and Steering Vectors
For a far-field direction parameterized by an elevation-azimuth pair $(\vartheta,\varphi)$ with elevation from the $xOy$ plane and azimuth from the $x$-axis, we define the departure direction vector as the unit propagation direction pointing from the array center to the far-field point
@@
Accordingly, we use the following indexed notational shorthands for distinct BS-user (BU), RIS-BS (RB), BS-RIS (BR), and RIS-user (RU) links, respectively, and define the direction vectors as follows
@@
$\!\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{RB}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{RB}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}\!,\!\!$ (7c)
$\!\boldsymbol{t}_{k,q}^{\text{RU}}(\boldsymbol{r}^{\text{R}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,q}^{\text{RU}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,q}^{\text{RU}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}\!.\!\!$ (7d)
@@
$\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{R}})=\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{RT}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{RT}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}.$ (9b)
@@ II-C Rotation-Aware Multipath Channels
$\boldsymbol{n}(\boldsymbol{r}^{\text{A}})=\boldsymbol{R}(\boldsymbol{r}^{\text{A}})\bar{\boldsymbol{n}},$ (10)
@@
where $G_{0}^{\text{B}}$ and $G_{0}^{\text{R}}$ denotes the maximum gain achieved when the signal aligns with the boresight, and $b^{\text{B}}$ and $b^{\text{R}}$ are the directivity factors controlling the main-lobe width, following from directivity normalization over the full sphere. Then, with given numbers of resolvable paths $L_{k}$, $P$, and $Q_{k}$, and small-scale complex gains $\delta_{k,l}$, $\gamma_{p}$, and $\psi_{k,q}$ that absorb path loss and fading, the BS-to-user channel $\boldsymbol{h}_{k}\left(\boldsymbol{r}^{\text{B}}\right)\in\mathbb{C}^{M}$, BS-to-RIS channel $\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\in\mathbb{C}^{M\times N}$, and RIS-to-user channel $\boldsymbol{g}_{k}\left(\boldsymbol{r}^{\text{R}}\right)\in\mathbb{C}^{N}$ are respectively expressed as
$\boldsymbol{h}_{k}\left(\boldsymbol{r}^{\text{B}}\right)=\sum_{l=1}^{L_{k}}{\delta_{k,l}\sqrt{G_{k,l}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}}\boldsymbol{t}_{k,l}^{\text{BU}}\left(\boldsymbol{r}^{\text{B}}\right),$ (12a)
$\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)=\sum_{p=1}^{P}{\gamma_{p}\sqrt{G_{p}^{\text{B}}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)G_{p}^{\text{R}}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}}$
$\qquad\qquad\qquad\quad\,\,\times\boldsymbol{t}_{p}^{\text{BR}}\left(\boldsymbol{r}^{\text{B}}\right)\left(\boldsymbol{t}_{p}^{\text{RB}}\left(\boldsymbol{r}^{\text{R}}\right)\right)^{H},$ (12b)
$\boldsymbol{g}_{k}\left(\boldsymbol{r}^{\text{R}}\right)=\sum_{q=1}^{Q_{k}}{\psi_{k,q}\sqrt{G_{k,q}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}}\boldsymbol{t}_{k,q}^{\text{RU}}\left(\boldsymbol{r}^{\text{R}}\right).$ (12c)
@@
$\boldsymbol{f}_{k}(\boldsymbol{r}^{\text{B}}\!,\boldsymbol{r}^{\text{R}}\!,\boldsymbol{\theta})\!=\!\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})\!+\!\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\mathrm{diag}(\boldsymbol{\theta})\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}}),\!\!$ (13)
@@
$\boldsymbol{f}_{\text{S},a}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})=$ $\sqrt{G_{a}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}\boldsymbol{t}_{a}^{\text{BT}}\left(\boldsymbol{r}^{\text{B}}\right)+\!\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\mathrm{diag}(\boldsymbol{\theta})$
$\times\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{R}}).$ (14)
II-D Transmit Signal, Communication SINR, and Sensing NMSE
@@
where $\boldsymbol{W}_{\text{C}}=[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{K}\boldsymbol{]}\in\mathbb{C}^{M\times K}$ collects $K$ communication beams, $\boldsymbol{W}_{\text{S}}\in\mathbb{C}^{M\times M}$ contains dedicated sensing beams, $\boldsymbol{s}_{\text{C}}=[s_{\text{C},1},\dots,s_{\text{C},K}]^{T}$ denotes the communication symbols, and $\boldsymbol{s}_{\text{S}}$ denotes the sensing signals, with $\mathbb{E}\left[\boldsymbol{s}_{\text{S}}\boldsymbol{s}_{\text{S}}^{H}\right]=\boldsymbol{I}_{M}$, $\boldsymbol{s}_{\text{C}}\sim\mathcal{C}\mathcal{N}\left(\boldsymbol{0},\boldsymbol{I}_{K}\right)$, and $\mathbb{E}\left[\boldsymbol{s}_{\text{S}}\boldsymbol{s}_{\text{C}}^{H}\right]=0$. Let $\boldsymbol{W}\triangleq\left[\boldsymbol{W}_{\text{C}},\boldsymbol{W}_{\text{S}}\right]$ denotes the BS ISAC transmit beamforming matrix, $\boldsymbol{s}\triangleq\left[\boldsymbol{s}_{\text{C}}^{T},\boldsymbol{s}_{\text{S}}^{T}\right]^{T}$, the received signal at the communication user $k$ is written as
$c_{k}$ $=\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{Ws}$ (16a)
$=\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{w}_{k}s_{\text{C},k}+\sum_{i=1,i\neq k}^{K}{\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{w}_{i}}s_{\text{C},i}$
@@
where $n_{k}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{k}^{2}\right)$ is the complex additive white Gaussian noise. Taking further $\boldsymbol{W}\triangleq\left[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{K+M}\right]$, where $\boldsymbol{w}_{i}$ is the $i$-th column of $\boldsymbol{W}$, the SINR of the $k$-th user is
@@
Given a discretized grid $\{(\vartheta_{a},\varphi_{a})\}_{a=1}^{A},a\in\{1,\dots,A\}$ with $A$ sampled points in total, which samples the azimuth and elevation angular domain over the target sensing area $\mathcal{A}$, along with the desired beampattern $\mathcal{P}_{\text{d}}(\vartheta_{a},\varphi_{a})$, the mean squared error (MSE) between $\mathcal{P}$ and $\mathcal{P}_{\text{d}}$ to evaluate the similarity of the sensing beampattern is defined as [41]
@@
A typical spotlight-type desired beampattern is given by
@@
where $\mathcal{F}$ is the set of focused grid points in the region $\mathcal{A}$. While MSE is intuitive, it is scale-sensitive: the same beampattern shape may lead to very different MSE values when the overall beampattern level fluctuates due to transmit-power changes, propagation loss, blockage, or orientation-dependent element gains, which are prevalent in low-altitude geometries. To obtain a dimensionless and scale-robust metric that emphasizes shape fidelity rather than absolute amplitude, we adopt the NMSE metric with a scaling parameter $\iota>0$ as
$\text{NMSE}\triangleq\frac{\sum_{a=1}^{A}{\left|\mathcal{P}_{a}-\iota\mathcal{P}_{d,a}\right|^{2}}}{\sum_{a=1}^{A}{\left|\iota\mathcal{P}_{d,a}\right|^{2}}}=\frac{\left\|\boldsymbol{p}-\iota\boldsymbol{p}_{\text{d}}\right\|_{2}^{2}}{\left\|\iota\boldsymbol{p}_{\text{d}}\right\|_{2}^{2}}.$ (22)
where $\boldsymbol{p}=[\mathcal{P}_{1},\ldots,\mathcal{P}_{A}]^{T}$ and $\boldsymbol{p}_{\text{d}}=[\mathcal{P}_{\text{d},1},\ldots,\mathcal{P}_{\text{d},A}]^{T}$. The scaling parameter $\iota$ decouples the global amplitude mismatch from the spatial shape mismatch: even when $\mathcal{P}_{a}$ is uniformly scaled by channel loss or array gain variations, $\text{NMSE}(\iota)$ can remain small as long as the angular shape matches. Moreover, using $\|\iota\boldsymbol{p}_{\text{d}}\|_{2}^{2}$ in the denominator avoids the trivial solution $\iota=0$, which would artificially reduce the numerator, thereby enforcing a meaningful normalization [^2] The adopted beampattern matching NMSE is a design-friendly proxy for sensing quality: a smaller NMSE indicates better mainlobe fidelity in the intended angular sectors and reduced leakage or sidelobes elsewhere, which typically improves the received peak-to-maximum-sidelobe ratio (PSLR) and signal-to-clutter-plus-noise ratio (SCNR), and the reliability of detection or estimation in the region of interest.
For fixed $(\boldsymbol{W},\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, we update $\iota$ by minimizing $\text{NMSE}(\iota)$ in closed form. Let $c=\|\boldsymbol{p}\|_{2}^{2}$, $d=\boldsymbol{p}_{\text{d}}^{T}\boldsymbol{p}$, and $e=\|\boldsymbol{p}_{\text{d}}\|_{2}^{2}$. When $d>0$, the minimizer is
$\iota^{\star}=\arg\min_{\iota>0}\text{NMSE}(\iota)=\frac{\|\boldsymbol{p}\|_{2}^{2}}{\boldsymbol{p}_{\text{d}}^{T}\boldsymbol{p}}.$ (23)
In practice, to avoid numerical issues when $\boldsymbol{p}_{\text{d}}^{T}\boldsymbol{p}$ is very small (e.g., at initialization), we use a safeguarded update $\iota^{\star}=\|\boldsymbol{p}\|_{2}^{2}/(\boldsymbol{p}_{\text{d}}^{T}\boldsymbol{p}+\varepsilon)$ with a small $\varepsilon>0$. With $\iota=\iota^{\star}$, the NMSE in (22) measures how well the synthesized pattern matches the desired pattern up to an overall scaling, which is particularly suitable when (i) the absolute beampattern level is affected by the BS power budget and propagation loss, (ii) the sensing objective primarily concerns mainlobe or sidelobe shape rather than absolute amplitude, and (iii) the system expects the sensing metric to be comparable across different system settings, e.g., $P_{\text{B}}$, $M$, NN, and user or target geometries.
@@ II-E Problem Formulation
Accordingly, we jointly optimize the BS and RIS rotations, the RIS phase shifts, and the BS transmit beamforming to maximize a weighted-sum utility of downlink communication sum rate and sensing beampattern fidelity. Specifically, we consider
$\text{(P1)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\right)}-\rho\text{NMSE}$ (24a)
s.t. $\|\boldsymbol{W}\|_{F}^{2}=\mathrm{Tr}(\boldsymbol{WW}^{H})\leq P_{\text{B}},$ (24b)
$|\theta_{n}|=1,\forall n\in\mathcal{N},$ (24c)
$r_{\mathrm{i}}^{\text{B}}\!\in\left[r_{\mathrm{i},\min}^{\text{B}},r_{\mathrm{i},\max}^{\text{B}}\right],\mathrm{i}\in\{\mathrm{x},\mathrm{y},\mathrm{z}\},$ (24d)
$r_{\mathrm{i}}^{\text{R}}\!\in\left[r_{\mathrm{i},\min}^{\text{R}},r_{\mathrm{i},\max}^{\text{R}}\right],\mathrm{i}\in\{\mathrm{x},\mathrm{y},\mathrm{z}\},$ (24e)
where $\rho\geq 0$ in (24a) is a trade-off weight controlling the relative emphasis on communication throughput versus sensing beampattern matching, $P_{\text{B}}$ in (24b) is the BS power budget, the unit-modulus constraint (24c) enforces a passive RIS, and the box constraints (24d) and (24e) capture mechanical limits of the array rotations at BS and RIS, respectively. Note that the proposed formulation in (24a) directly targets the fundamental ISAC trade-off by optimizing a scalarized Pareto objective.
Compared with a strict threshold form of constraint on NMSE, the weighted-sum design offers three practical advantages. First, in low-altitude scenarios, the sensing pattern requirement may become temporarily difficult or even infeasible under fast geometry variations and strong multiuser interference, while the weighted-sum objective avoids infeasibility, being robust to geometry variation, and yields a graceful performance trade-off. Second, hard-threshold formulations typically require penalty or dual updates and may introduce non-smooth hinge-type terms, which can complicate algorithm design and slow convergence, while a weighted sum leads to a single-level smooth objective, enabling simpler tuning and more stable optimization. Third, by sweeping $\rho$, one can efficiently explore the Pareto frontier between the sum rate and NMSE, which provides a comprehensive characterization of the communication-sensing trade-off beyond enforcing a single scenario-dependent threshold.
From the formulated problem, we observe that the rotation variables $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ provide a geometric control that affects the steering vectors and directive gain, thereby complementing the electromagnetic control offered by active and passive beamforming at the RIS and BS arrays; consequently, jointly optimizing these blocks yields a significant DoF for low-altitude ISAC. However, (P1) is highly non-convex and challenging to solve directly due to unit-modulus constraints, rotation-induced nonlinearities in the steering vector and directive gain, and the coupling among $\boldsymbol{r}^{\text{B}}$, $\boldsymbol{r}^{\text{R}}$, $\boldsymbol{\theta}$, and $\boldsymbol{W}$ in both SINR and MSE expressions.